OpenStudy (anonymous):

compute the derivative where f(x)cos^(-1)(x) and x=1/2

OpenStudy (anonymous):

Is this the function?\[f(x)=\cos^{-1}x \] If so, then you just have to commit the derivative to memory (I can prove it if you want, but it's not needed). The derivative is: \[f'(x)=-1/\sqrt{1-x^2}\] Just plug x = 1/2 into f'(x) and you'll have your answer.

OpenStudy (anonymous):

yes thats the function.. just plug in 1/2?

OpenStudy (anonymous):

yup